# Distributions and Direct Parametrization for Stable Stochastic State-Space Models

### v1

@@ IV-B Distributions for SSMs satisfying the stochastic BRL
with $P,Q\in\mathbb{S}^{n}_{>}$, $S\in\mathbb{S}^{n}_{\mathrm{skew}}$, $\tilde{F}\in\mathbb{R}^{n\times n}$, $C\in\mathbb{R}^{q\times n}$, and $D\in\mathbb{R}^{q\times\ell},~{}\tilde{B},\tilde{G}\in\mathbb{R}^{n\times\ell}$ satisfying
$\tilde{B}^{\top}\tilde{B}+D^{\top}D+\tilde{G}^{\top}\tilde{G}<\gamma^{2}I.$ (10)
Conversely, any choice of $P,Q\in\mathbb{S}^{n}_{>}$, $S\in\mathbb{S}^{n}_{\mathrm{skew}}$, $\tilde{F}\in\mathbb{R}^{n\times n}$, $C\in\mathbb{R}^{q\times n}$, and $D\in\mathbb{R}^{q\times\ell},~{}\tilde{B},\tilde{G}\in\mathbb{R}^{n\times\ell}$ satisfying condition (10) with $A,G,B$ defined as in (9) will yield an SSM satefying the stochastic BRL with $\|\mathbb{L}\|<\gamma$.
@@
$M<0\Leftrightarrow M_{11}<0,~{}M_{22}-M_{12}M^{-1}_{11}M^{\top}_{12}<0.$ (11)
The inequality $M_{11}<0$ implies that there exists a $Q\in\mathbb{S}^{n}_{>}$ such that $PA+A^{\top}P+F^{\top}PF=-Q-C^{\top}C$. By defining the positive definite matrix $\tilde{Q}:=Q+C^{\top}C$, the equation becomes $PA+A^{\top}P+F^{\top}PF=\tilde{Q}$ which is a Lyapunov equation for internal stability (3) (replacing $\tilde{Q}$ with $Q$). Therefore, using Proposition 2, we can write $A=-\frac{1}{2}P^{-1}(Q+C^{\top}C+\tilde{F}^{\top}\tilde{F}+S)$ with $\tilde{F}\in\mathbb{R}^{n\times n}$ and $S\in\mathbb{S}^{n}_{\mathrm{skew}}$, and $F=L_{P^{-1}}\tilde{F}$.
We now show that for any SSM with matrices satisfying (5), we can find $\tilde{B},\tilde{G}$ with $\tilde{B}^{\top}\tilde{B}+D^{\top}D+\tilde{G}^{\top}\tilde{G}<\gamma^{2}I$. Using (11), we know that $M_{22}-M_{12}M_{11}^{-1}M_{12}^{\top}<0$ and $M_{11}=PA+A^{\top}P+F^{\top}PF+C^{\top}C:=-Q<0$ (internal stability). Now let $L_{Q^{-1}},L_{P}$ be the lower-triangular Cholesky factor for $Q^{-1}$ and $P$, respectively, $\tilde{B}=L^{-1}_{Q}\left(PB+\rho\tilde{F}^{\top}\tilde{G}+C^{\top}D\right)$ and $\tilde{G}=L^{-1}_{P}G$. Then (9b) follows, and by substiuting $\tilde{F},\tilde{G},\tilde{B}$ in $M_{11}-M_{12}M^{-1}_{11}M_{12}^{\top}$, we get the condition $\tilde{B}^{\top}\tilde{B}+D^{\top}D+\tilde{G}^{\top}\tilde{G}<\gamma^{2}I$. For the converse, direct substitution of $A,G,B$ defined in (9) with any $P,Q\in\mathbb{S}^{n}_{>}$, $S\in\mathbb{S}^{n}_{\mathrm{skew}}$, $\tilde{F}\in\mathbb{R}^{n\times n}$, $C\in\mathbb{R}^{q\times n}$, and $D\in\mathbb{R}^{q\times\ell},~{}\tilde{B},\tilde{G}\in\mathbb{R}^{n\times\ell}$ satisfying the condition (10), will produce a matrix $M$ as in (5) satisfying $M<0$. ∎
@@
$Z^{\top}Z<I,\quad\text{with}\quad Z:=\frac{1}{\gamma}\begin{bmatrix}\tilde{B}^{\top}~{}D^{\top}~{}\tilde{G}^{\top}\end{bmatrix}^{\top}.$ (12)
@@
To address this issue, we instead aim to sample $Z$ so that its maximum singular value $\sigma^{z}_{1}$ is close to 1. Specifically, we let $\sigma^{z}_{1}=(1-\varepsilon)$ with $10<\varepsilon<<$ 1. To implement this, we consider the Singular Value Decomposition (SVD) of $Z$:
@@
[…]
$\mathrm{vec}(\tilde{U}_{z})\sim\mathcal{N}(0,I\otimes\Sigma_{U_{z}})\quad\text{and}\quad\mathrm{vec}(\tilde{V}_{z})\sim\mathcal{N}(0,I\otimes\Sigma_{V_{z}}),$
and then compute $U_{z}$ and $V_{z}$ as the orthogonal factors in the QR-factorizations of $\tilde{U}_{z}$ and $\tilde{V}_{z}$, respectively. The resulting distribution over $U_{z}$ and $V_{z}$ is known as the matrix Angular Central Gaussian (ACG) distribution [16]. Another approach to sample $U_{z}$ and $V_{z}$ is to sample them using Cayley transform [17] $U_{z}=E_{u}(I-S_{u})(I+S_{u})^{-1}$ and $V_{z}=E_{z}(I-S_{v})(I+S_{v})^{-1}$, where $\mathrm{veck}(S_{u})\sim\mathcal{N}(\mu_{u},\Sigma_{u})$, $\mathrm{veck}(S_{v})\sim\mathcal{N}(\mu_{v},\Sigma_{v})$, and $E_{u},E_{v}$ are diagonal matrices with entries ±$\pm 1$ sampled uniformly.
@@
If we want to have more flexibility in specifying the uncertainty in the individual norms of $\tilde{B}$, $\tilde{G}$, and $D$, we sample them based on their SVDs (similar to $Z$) with the maximum singular values set to $\lambda_{b}(1-\varepsilon)\gamma$, $\lambda_{g}(1-\varepsilon)\gamma$, and $\lambda_{d}(1-\varepsilon)\gamma$, respectively. Here, the weights $\lambda_{b}$, $\lambda_{g}$, and $\lambda_{d}$ are strictly positive and satisfy $\lambda_{b}+\lambda_{g}+\lambda_{d}=1$.
These weights are drawn from a Dirichlet distribution: $\lambda=\left[\lambda_{b},\lambda_{g},\lambda_{d}\right]\sim\mathrm{Dir}(\alpha_{1},\alpha_{2},\alpha_{3}),$ with $\alpha_{1},\alpha_{2},\alpha_{3}>0$.
@@
Let $\gamma>0$, $P^{-1}\sim\mathrm{Wishart}(k_{p},\Sigma_{p})$ with $k_{p}>n$, $\mathrm{vec}(\tilde{F})\sim\mathcal{N}(\mu_{f},\Sigma_{f})$; $\mathrm{vec}(C)\sim\mathcal{N}(\mu_{c},\Sigma_{c})$, and $\mathrm{veck}(S)\sim\mathcal{N}(\mu_{s},\Sigma_{s})$. Let $Z=U_{z}\,\mathrm{diag}_{2n+q,\ell}(\sigma^{z}_{1},\dots,\sigma^{z}_{k})V^{\top}_{z},$ with $\sigma^{z}_{1}=1-\varepsilon$ for a fixed constant $0<\varepsilon\ll 1$, and the remaining singular values $\sigma^{z}_{2},\dots,\sigma_{k}^{z}$ are drawn independently from $\mathrm{U}(0,1-\varepsilon)$ with $k=\min(2n+q,\ell)$. The orthonormal matrices $U_{z}$ and $V_{z}$ are sampled from a matrix ACG distribution (or via a Cayley transform based on Gaussian skew-symmetric matrices and diagonal matrices with ±$\pm 1$ entries sampled uniformly), and the matrices $\tilde{B}$, $D$, and $\tilde{G}$ are specified as in (12). If all the random variable are independent with $A,B,C,D,F,G$ defined via (9), then resulting SSM in (1) is said to follow
  • The WNS-BRL-ACG (WNS-BRL-Cayley) distribution if $Q\in\mathbb{S}^{n}_{>}$ is fixed and $U_{z},V_{z}$ follow the ACG (or Cayley) distribution.
@@ V Bayesian Inference Example
In this section, we illustrate how the prior distributions for the stable SSMs developed in Sections IV-A and subsequent subsections can be employed in Bayesian inference. In our setting, we assume that the true SSM (1) is known to satisfy the stochastic BRL with $\|\mathbb{L}\|<\gamma$, where $\gamma>0$ may itself be modeled as a random variable to capture uncertainty in the system gain.
V-A Bayesian Inference Setup
Let $\{t_{i}\}_{i=1}^{N}\subset[0,T]$ denote the measurement time instances. At each $t_{i}$, we record the input $u(t_{i})$ and the corresponding noisy output $\tilde{y}(t_{i})$, where $\tilde{y}(t_{i})=y(t_{i};x_{0},u)+\xi(t_{i}).$ Here, $y(t;x_{0},u)$ is the output of the SSM (1) corresponding to the initial condition $x(0)=x_{0}\in\mathbb{R}^{n}$, and $\xi(t_{i})$ represents measurement noise. We assume that the noise $\xi(t_{i})$ is independent and identically distributed $\xi(t_{i})\sim\mathcal{N}(0,\sigma^{2}I_{q})$. Furthermore, we assume that the data is collected from $M$ independent realizations (sample paths) of the SSM, all starting from the same initial condition $x_{0}$. With this setup, Bayesian inference can be performed to update the prior distributions over the SSM parameters based on the observed data $\{(u(t_{i}),\tilde{y}(t_{i}))\}_{i=1}^{N}$ from the $M$ experiments.
To generate data, we sampled a model with $n=2,q=1,\ell=2$ using the $Q$-WNS-BRL-Cayley distribution with $\Sigma_{p}=I_{n},\Sigma_{q}=2I,k_{p}=6,k_{q}=6,\mu_{s}=0,\Sigma_{s}=0.01I,\mu_{f}=0,\Sigma_{f}=2I,\mu_{c}=0,\Sigma_{c}=2I,\varepsilon=10^{-4},\Sigma_{u}=I$, and $\Sigma_{v}=I$. We also set $\rho=0.3$, $\gamma=3$, and the measurement error variance $\sigma=0.15$. We took $M=3$ experiments and generated the inputs using 6 Fourier terms with random coefficients.
The priors used for the inference were chosen to be different from the ones used for sampling the model to generate the data. Specifically, we choose the WNS-BRL-Cayley with fixed $Q=I$ and with the other different choices being $k_{p}=4,k_{q}=4$, and $\Sigma_{s}=I$. In addition, We assigned a prior $\rho\sim\mathrm{U}(-0.95,0.95)$ and $\gamma\sim\mathrm{Gamma}(1,1)$ ($\mathbb{E}[\gamma]=1$).
To perform the inference, we used a Markov Chain Monte Carlo (MCMC) [18] method to sample from the posterior distribution given the data. The MCMC method we used is Hamiltonian Monte Carlo (HMC) with a step size of 0.01 and a number of steps being 3 [19]. We implemented the inference using the Turing.jl package in Julia [20]. The code used for generating the results is available at Gitlab https://bit.ly/4hcKnQs.
To compare, we used HMC with the same settings but with priors assigned directly on the SSM matrices without any parametrization. This approach can cause numerical issues during an MCMC sampling strategy since sampling an unstable SSM is plausible. To deal with this issue, every time we detect an unstable trajectory in the MCMC sampling strategy, we assign a log-likelihood of $\infty$- ∞ to that particular sample of matrices. Fig. 1 and Fig. 2 demonstrate the results. In Fig. 1, we compare the posterior mean and 1000 posterior trajectory samples of the mean output for the two prior modeling strategies. In each case, we generated the results by first drawing a posterior sample of the system matrices, then simulating 100 trajectories of the SSM in (1) with those matrices, and finally computing their sample mean. These mean trajectories are overlaid for 1000 posterior samples (in blue), along with the true mean trajectory (in black) which was generated using the actual model parameters.
As seen in the top panel of Fig.1, the trajectories obtained under the WNS-BRL-Cayley prior remain close to the true mean throughout the entire inference and prediction horizon. Even after the inference period ends (delineated by the green dashed line), the simulated trajectories exhibit bounded and stable behavior. This observation highlights that incorporating stability-inducing priors—here, through the WNS-BRL-Cayley formulation—naturally prevents the inferred system from diverging when extrapolated beyond the observed data.
In contrast, the trajectories obtained under the free parametrization begin to diverge from the true mean shortly after the inference interval. Although we mitigate the instabilities by assigning a log-likelihood of $\infty$- ∞ to any sample that yields an unstable trajectory during the inference with the data, that does not ensure the posterior draws of the matrices of the SSM in (1) lead to an externally stable system.
A similar phenomenon is observed in Fig.2, which shows the variance of the output trajectories computed in the same way as in Fig.1. Under the WNS-BRL-ACG prior, the posterior trajectory samples remain well-behaved, producing a variance that closely matches the true variance over time. In contrast, the free parametrization exhibits a rapidly increasing variance once we move past the inference interval. These diverging variance trajectories again reflect the fact that we cannot guarantee with a free parametrization approach that the SSM in (1) is externally stable.
Overall, these figures illustrate that enforcing stability through the WNS-BRL-Cayley distribution yields posterior inferences that are robust and predictive performance that remains consistent even in the extrapolation regime. By contrast, a naive free parametrization of the SSM can cause significant numerical challenges during MCMC sampling and lead to unreliable predictions, particularly for longer-term forecasts.
Figure 1: The posterior mean (in red) and the posterior samples (in blue) of the mean over 100 realizations of the SSM in (1) compared to the mean over 100 realizations from the true SSM (in black).
@@ V-A Bayesian Inference Setup
Figure 2: The posterior mean (in red) and the posterior samples (in blue) of the variance over 100 realizations of the SSM in (1) compared to the variance over 100 realizations from the true SSM (in black).

### v2

@@ IV-B Distributions for SSMs satisfying the stochastic BRL
with $P,Q\in\mathbb{S}^{n}_{>0}$, $S\in\mathbb{S}^{n}_{\mathrm{skew}}$, $\tilde{F}\in\mathbb{R}^{n\times n}$, $C\in\mathbb{R}^{q\times n}$, $D\in\mathbb{R}^{q\times\ell},~\tilde{B},\tilde{G}\in\mathbb{R}^{n\times\ell}$ satisfying
$\tilde{B}^{\top}\tilde{B}+D^{\top}D+\tilde{G}^{\top}\tilde{G}<\gamma^{2}I,$ (11)
and $L_{P^{-1}},L_{Q}$ are the lower-triangular Cholesky factors of $P^{-1}$ and $Q$, respectively. Conversely, any choice of $P,Q\in\mathbb{S}^{n}_{>0}$, $S\in\mathbb{S}^{n}_{\mathrm{skew}}$, $\tilde{F}\in\mathbb{R}^{n\times n}$, $C\in\mathbb{R}^{q\times n}$, and $D\in\mathbb{R}^{q\times\ell},~\tilde{B},\tilde{G}\in\mathbb{R}^{n\times\ell}$ satisfying condition (11) with $A,F,G,B$ defined as in (10) will yield an SSM satisfying the stochastic BRL with $\|\mathbb{L}\|<\gamma$.
@@
$M<0\Leftrightarrow M_{11}<0,~M_{22}-M^{\top}_{12}M^{-1}_{11}M_{12}<0.$ (12)
Now define $Q\in\mathbb{S}^{n}_{>0}$ by $Q=M_{11}=PA+A^{\top}P+F^{\top}PF+C^{\top}C$. Then $PA+A^{\top}P+F^{\top}PF=-(Q+C^{\top}C)$ is a Lyapunov equation for internal stability (3). Therefore, using Proposition 2, we can write $A=-\frac{1}{2}P^{-1}(Q+C^{\top}C+\tilde{F}^{\top}\tilde{F}+S)$ with $\tilde{F}\in\mathbb{R}^{n\times n}$, $S\in\mathbb{S}^{n}_{\mathrm{skew}}$, and $F=L_{P^{-1}}\tilde{F}$.
We now show that for any SSM with matrices satisfying (5), we can find $\tilde{B},\tilde{G}$ with $\tilde{B}^{\top}\tilde{B}+D^{\top}D+\tilde{G}^{\top}\tilde{G}<\gamma^{2}I$. Using (12), we know that $M_{22}-M^{\top}_{12}M_{11}^{-1}M_{12}<0$ and $M_{11}\!=\!-Q\!<\!0$. Now let $\tilde{B}=L^{-1}_{Q}\left(PB+\rho\tilde{F}^{\top}\tilde{G}+C^{\top}D\right)$ and $\tilde{G}=L^{-1}_{P^{-1}}G$, then (10b) follows, and by substituting $\tilde{F},\tilde{G},\tilde{B}$ in $M_{22}-M_{12}^{\top}M^{-1}_{11}M_{12}$, we get the condition (11). For the converse, direct substitution of $A,F,G,B$ defined in (10) with any $P,Q\in\mathbb{S}^{n}_{>0}$, $S\in\mathbb{S}^{n}_{\mathrm{skew}}$, $\tilde{F}\in\mathbb{R}^{n\times n}$, $C\in\mathbb{R}^{q\times n}$, and $D\in\mathbb{R}^{q\times\ell},~\tilde{B},\tilde{G}\in\mathbb{R}^{n\times\ell}$ satisfying the condition (11), will produce a matrix $M$ as in (5) satisfying $M<0$. ∎
Note that the $Q$ in Proposition 3 is a positive‑definite slack matrix that is different from the $Q$ in Proposition 2.
@@
$Z^{\top}Z<I,\quad\text{with}\quad Z\mathrel{\mathop{\ordinarycolon}}=\frac{1}{\gamma}\begin{bmatrix}\tilde{B}^{\top}~D^{\top}~\tilde{G}^{\top}\end{bmatrix}^{\top}$. (13)
@@
To address this issue, we instead aim to sample $Z$ so that its maximum singular value $\sigma^{z}_{1}$ is close to 1. Specifically, we let $\sigma^{z}_{1}=(1-\varepsilon)$ with $0<\varepsilon<<1$. To implement this, we consider the Singular Value Decomposition (SVD) of $Z$:
@@
with the maximum singular value $\sigma^{z}_{1}$ fixed to $(1-\varepsilon)$ and the rest are independently sampled from the uniform distribution $\sigma^{z}_{i}\sim\mathrm{U}\bigl(0,\,1-\varepsilon\bigr),$ for $i\in\{2\dots,k\}$. For the orthonormal matrices $U_{z}$ and $V_{z}$, one practical approach is to sample $\mathrm{vec}(\tilde{U}_{z})\sim\mathcal{N}(0,I\otimes\Sigma_{U_{z}})\text{ and }\mathrm{vec}(\tilde{V}_{z})\sim\mathcal{N}(0,I\otimes\Sigma_{V_{z}}),$ and then compute $U_{z}$ and $V_{z}$ as the orthogonal factors in the QR-factorizations of $\tilde{U}_{z}$ and $\tilde{V}_{z}$, respectively. The resulting distribution over $U_{z}$ and $V_{z}$ is known as the matrix Angular Central Gaussian (ACG) distribution [21]. Another approach to sample $U_{z}$ and $V_{z}$ is to sample them using Cayley transform [22] $U_{z}\!=\!E_{u}(I-S_{u})(I+S_{u})^{-1}$ and $V_{z}\!=\!E_{z}(I-S_{v})(I+S_{v})^{-1}$, where $\mathrm{veck}(S_{u})\sim\mathcal{N}(\mu_{u},\Sigma_{u})$, $\mathrm{veck}(S_{v})\sim\mathcal{N}(\mu_{v},\Sigma_{v})$, and $E_{u},E_{v}$ are diagonal matrices with entries $\pm 1$ sampled uniformly.
@@
If we want to have more flexibility in specifying the uncertainty in the individual norms of $\tilde{B}$, $\tilde{G}$, and $D$, we sample them based on their SVDs (similar to $Z$) with the maximum singular values set to $\lambda_{b}(1-\varepsilon)\gamma$, $\lambda_{g}(1-\varepsilon)\gamma$, and $\lambda_{d}(1-\varepsilon)\gamma$, respectively. Here, the weights $\lambda_{b}$, $\lambda_{g}$, and $\lambda_{d}$ are strictly positive and satisfy $\lambda_{b}+\lambda_{g}+\lambda_{d}=1$. These weights can be drawn from a Dirichlet distribution $\lambda=\left[\lambda_{b},\lambda_{g},\lambda_{d}\right]\sim\mathrm{Dir}(\alpha_{1},\alpha_{2},\alpha_{3}),$ with $\alpha_{1},\alpha_{2},\alpha_{3}>0$.
@@
Let $\gamma\!>\!0$, $P^{-1}\!\sim\!\mathrm{Wishart}(k_{p},\Sigma_{p})$ with $k_{p}\!\geq\!n$, $\mathrm{vec}(\tilde{F})\!\sim\!\mathcal{N}(\mu_{f},\Sigma_{f})$; $\mathrm{vec}(C)\sim\mathcal{N}(\mu_{c},\Sigma_{c})$, and $\mathrm{veck}(S)\!\sim\!\mathcal{N}(\mu_{s},\Sigma_{s})$. Let $Z=U_{z}\,\mathrm{diag}_{2n+q,\ell}(\sigma^{z}_{1},\dots,\sigma^{z}_{k})V^{\top}_{z},$ with $\sigma^{z}_{1}=1-\varepsilon$ for a fixed constant $0<\varepsilon\ll 1$, and the remaining singular values $\sigma^{z}_{2},\dots,\sigma_{k}^{z}$ are drawn independently from $\mathrm{U}(0,1-\varepsilon)$ with $k=\min(2n+q,\ell)$. The orthonormal matrices $U_{z}$ and $V_{z}$ are sampled from a matrix ACG distribution (or via a Cayley transform based on Gaussian skew-symmetric matrices and diagonal matrices with $\pm 1$ entries sampled uniformly), and the matrices $\tilde{B}$, $D$, and $\tilde{G}$ are specified as in (13). If all the random variable are independent with $A,B,C,D,F,G$ defined via (10), then resulting SSM in (1) is said to follow
  • The WNS-BRL-ACG (WNS-BRL-Cayley) distribution if $Q\in\mathbb{S}^{n}_{>0}$ is fixed and $U_{z},V_{z}$ follow the ACG (or Cayley) distribution.
@@
Remark 2
Even with no prior knowledge of the SSM, Proposition 3 provides an unconstrained parametrization for maximum likelihood estimation without the need for distributions. Similar to [15], positive scalars can be parametrized with an exponential function (e.g., $\alpha=\mathrm{e}^{\bar{\alpha}},\bar{\alpha}\in\mathbb{R}$) and bounded scalars by a sigmoid function (e.g., $\sigma_{z}^{i}=\frac{1-\varepsilon}{1+\mathrm{e}^{-{\bar{\sigma}^{i}_{z}}}},\bar{\sigma}_{z}^{i}\in\mathbb{R}$); positive-definite matrices via Cholesky factors postive diagonal entries; and orthogonal matrices through a Cayley transform. Alternatively, one may instead perform Riemannian gradient descent on the manifolds of positive-definite, skew-symmetric and orthogonal matrices, similar to [13].
@@ V Bayesian Inference Example
Figure 1: The posterior mean (in red) and the posterior samples (in blue) of the mean (variance) over 100 realizations of the SSM in (1) compared to the mean (variance) over 100 realizations from the true SSM (in black).
In this section, we illustrate how the prior distributions from Definition 4 can be employed in Bayesian inference. In our setting, we assume that the true SSM (1) is known to satisfy the stochastic BRL with $\|\mathbb{L}\|<\gamma$, where $\gamma>0$ may itself be modeled as a random variable to capture uncertainty in the system gain.
Let $\{t_{i}\}_{i=1}^{N}\subset[0,T]$ denote the measurement time instances. At each $t_{i}$, we record the input $u(t_{i})$ and the corresponding noisy output $\tilde{y}(t_{i})$, where $\tilde{y}(t_{i})=y(t_{i};x_{0},u)+\xi(t_{i}).$ Here, $y(t;x_{0},u)$ is the output of the SSM (1) corresponding to the initial condition $x(0)=x_{0}$, and $\xi(t_{i})$ represents measurement noise. We assume that the noise $\xi(t_{i})$ is independent and identically distributed $\xi(t_{i})\sim\mathcal{N}(0,\sigma^{2}I_{q})$. Furthermore, we assume that the data is collected from $M$ independent realizations (sample paths) of the SSM, all starting from the same initial condition $x_{0}$. With this setup, Bayesian inference can be performed to update the prior distributions over the SSM parameters based on the observed data $\{(u(t_{i}),\tilde{y}(t_{i}))\}_{i=1}^{N}$ from the $M$ experiments.
To generate data, we sampled a model with ${n=4},q=1,\ell=2$ using the $Q$-WNS-BRL-Cayley distribution with $\Sigma_{p}\!=\!I_{n},\Sigma_{q}\!=\!2I,k_{p}\!=\!6,k_{q}\!=\!6,\mu_{s}\!=\!0,\Sigma_{s}\!=\!0.01I,\mu_{f}\!=\!0,\Sigma_{f}\!=\!2I,\mu_{c}\!=\!0,\Sigma_{c}\!=\!2I,\varepsilon=10^{-4},\Sigma_{u}\!=\!I$, and $\Sigma_{v}\!=\!I$. We also set $\rho\!=\!0.3$, $\gamma\!=\!3$, and the measurement error variance $\sigma\!=\!0.15$. We took M=10M\!=\!10 experiments and generated the inputs using 6 Fourier terms with random coefficients. The priors used for the inference were chosen to be different from the ones used for sampling the model to generate the data. Specifically, we chose the $Q$-WNS-BRL-Cayley with the other different choices being $k_{p}=4,k_{q}=4,\Sigma_{q}=I$, and $\Sigma_{s}=I$. In addition, We assigned a prior $\rho\sim\mathrm{U}(-0.95,0.95)$ and $\gamma\sim\mathrm{Gamma}(1,1)$ ($\mathbb{E}[\gamma]=1$).
To perform the inference, we used a Markov Chain Monte Carlo (MCMC) [23] method to obtain samples from the posterior distribution given the data. The MCMC method we used is Hamiltonian Monte Carlo (HMC) with a step size of 0.01 and a number of steps being 3 [24]. We implemented the inference using the Turing.jl package in Julia [25]. The code used for generating the results is available at Github https://github.com/MOHAMMADZAHD93/StableSSMs.
To compare, we used HMC with the same settings but with priors assigned directly on the SSM matrices without any parametrization. This approach can cause numerical issues during the sampling process of an MCMC strategy since proposed SSM samples can lead to unstable trajectories. To deal with this, every time a proposed SSM sample results in an unstable trajectory, we assign a log-likelihood of -$\infty$ to that particular sample of matrices. That is, we force a rejection of the proposed sample in the MCMC step. Fig. 1 demonstrate the results. In Fig. 1a, we compare the posterior mean and 500 posterior trajectory samples of the mean output (computed over 100 sample paths) for the two prior modeling strategies. In each case, we generated the results by simulating 100 trajectories for each one of the last 1000 posterior SSM samples obtained by the HMC and computing their sample mean trajectory. The last 500 mean trajectories obtained from the last 500 SSM posterior samples out of the last 1000 samples are overlaid (in blue) along with the true mean trajectory (in black) which was generated using the actual model parameters.
As seen in the top panel of Fig.1a, the trajectories obtained under the $Q$-WNS-BRL-Cayley prior remain close to the true mean throughout the entire inference and prediction horizon. Even after the inference period ends (delineated by the green dashed line), the simulated trajectories exhibit bounded and stable behavior. This observation highlights that incorporating stability-inducing priors, done here through the $Q$-WNS-BRL-Cayley prior, naturally prevents the inferred system from diverging when extrapolated beyond the observed data. In contrast, the trajectories obtained under the free parametrization begin to diverge from the true mean shortly after the inference interval. Although we mitigate the instabilities by assigning a log-likelihood of -$\infty$ to any sample that yields an unstable trajectory during the MCMC sampling, that does not ensure the posterior draws of the matrices of the SSM in (1) lead to an externally stable system. A similar phenomenon is observed in Fig.1b, which shows the variance of the output trajectories computed in the same way as in Fig.1a. Under the $Q$-WNS-BRL-Cayley prior, the posterior trajectory samples remain well-behaved, producing a variance that closely matches the true variance over time. In contrast, the free parametrization exhibits a rapidly increasing variance once we move past the inference interval. These diverging variance trajectories again reflect the fact that we cannot guarantee with a free parametrization approach that the SSM in (1) is externally stable.
Overall, these figures illustrate that enforcing stability through the $Q$-WNS-BRL-Cayley distribution yields posterior samples that are robust and exhibit predictive performance that remains consistent even in the extrapolation regime. By contrast, a naive free parametrization of the SSM can cause numerical challenges during MCMC sampling and lead to unreliable predictions, particularly for longer-term forecasts.